A bag contains seven white and five black discs. Three disks are chosen from the bag. Find the probability so that after each draw the disc is replaced by one of the opposite colour.
Answer by VFBundy(438)

Probability of WHITE-BLACK-WHITE = 7/12 * 5/11 * 6/10 = 210/1320

Probability of BLACK-WHITE-BLACK = 5/12 * 7/11 * 4/10 = 140/1320

Probability of either = 210/1320 + 140/1320 = 350/1320 = 35/132